Grade 11Physical Chemistry

In an aqueous solution at 25 degree celsius has twice as many OH- ions as pure water its pOH is?

MA10
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
)]-7)and we know that at 25 degree C, pH+pOH=14,so calculate pH=14-pOH=14-[-log(2*10-7Mso pOH=-log(2*10-7 MAs per the questuion, OH-=2*10-7The hydroxide ions for pure water at 25C=10